# Mass-loaded magnetic explosions in the context of Magnetar Giant Flares and Fast Radio Bursts

### v1

@@ 4.2.2 Mass
The density profile of the plasmoid is related to the flux function $\Psi$, equations (17) and (21) multiplied by the prefactor $\Gamma^{3}v^{3}/r^{3}$ as demonstrated in equation (15). The two families of solutions we have studied here, for positive and negative $c_{1}$ have qualitatively different density profiles. For $c_{1}>0$, the minimum is located at the centre of the sphere ($r=0$) and peaks close to the location of the maximum of $\Psi$ with some modulation due to the $\Gamma^{3}$ prefactor that will push the maximum closer to the surface. On the contrary for $c_{1}<0$, the minimum is located close to the maximum of $\Psi$.
Let us first consider the rest mass contained within the expanding plasmoid. We integrate the rest mass density expression given by equation (15) within a sphere of given radius $R$, so that $R=cv_{0}\penalty 10000\ t$ at some time $t$, therefore corresponding to the physical location of the non-zero root of $g$. Accounting for the axial symmetry of the system, integration over $\phi$ provides a factor of $2\pi$.
@@
We confirm, as expected, that the rest mass contained within a spherical configuration does not evolve with time or the physical radius of the fluid, as the integral is on the $v$-space, where the solution does not evolve. However, the amount of mass contained inside the sphere, depends on the solution. In what follows we shall assume that the function $Q$ is constant and sets the scale for the mass. Furthermore, the total mass is sensitive to the choice of the additive constant, $\tilde{p}_{0}$. Here we make the following choices: for $c_{1}>0$, we set $\tilde{p}_{0}=0$, whereas for $c_{1}<0$ we choose the minimum acceptable value of $\tilde{p}_{0}$ so that the minimum of the density is exactly zero, thus $\tilde{p}_{0}=-\frac{c_{1}}{16\pi^{3}}$. The resulting masses from the integration are shown in Figure 7. We notice that for higher absolute values of $c_{1}$ the mass is higher. We note that the assumption of an additive constant affects the direct comparison of the mass: the P-type solutions have zero density on their surface, where the N-type ones have a non-zero density on their surfaces. To produce a reasonable comparison, if one considers the same density on the surface $v=v_{0}$ the P-type solutions have a higher mass load, compared to the N-type ones.
Figure 7: The mass $M$ and work $W$ for the P and N solutions as functions of $|c_{1}|$. Note that while $M$ and $W$ are plotted on the same axis the units differ. Moreover, $W$ evolves with time being inversely proportional to $t$.
@@ 4.2.3 Thermal Pressure Work
Following a similar procedure as in the density we can evaluate the work $W$ of the thermal pressure. It can be calculated through the integral of $pdV$. Upon integration for a sphere of radius $R$ at time $t$, we obtain the following expression:
@@
We notice that $W$ is inversely proportional to $t$, signifying the expansion of the system. This work is associated to the spatial redistribution of the plasma mass and pressure as the system expands. We further note, that the thermal pressure on the the surface of the sphere decreases with time, thus as the system expands less work in done against the external medium. The results of this integration for various choices is shown in Figure 7, with the dotted lines.
@@ 4.2.4 Temperature
Given the relation between the density and the pressure we can evaluate the temperature of the plasmoid, through the relation $T=\frac{m\penalty 10000\ p}{k_{B}\rho}$, where $m$ is the particle mass and $k_{B}$ is the Boltzmann constant. Therefore, the temperature is given by the form:
$T=\frac{m}{k_{B}}\frac{\tilde{p}}{\tilde{\rho}}\frac{\Gamma v}{r}=\frac{m}{k_{B}}\frac{\tilde{p}}{\tilde{\rho}}\frac{\Gamma}{ct}$ (27)
@@
As $\tilde{p}$ and $\tilde{\rho}$ are related to each other through $Q$, we are going to follow the assumption of a constant value of $Q(v,\theta)$ in the evaluation of the temperature, as we have already chosen in the calculation of mass through the density. Therefore, we expect a temperature profile for the spherical structure that decreases inversely proportionally with time.
For positive values of $c_{1}$ the maximum value of the temperature coincides approximately with the maximum of the flux function $\Psi$, while it is cooler at the boundaries of the sphere and the axis, Figure 4 panels c and g. For negative values of $c_{1}$ the minimum of the temperature occurs close to the maximum of $\Psi$, while the surroundings are hotter as illustrated in Figure 5 panels c and g. For the normalisation values assumed here, see the caption of Table LABEL:tab:1, and assuming a proton-electron plasma the units of the temperature correspond to $10^{11}$K.
@@ 5 Astrophysical implications
The solutions derived here provide a framework for the study of astrophysical systems of relativistically expanding configurations of magnetised plasma. It has been proposed that such structures may arise from magnetars expelling mass-loaded magnetic structures, which are associated to magnetar flares (Thompson & Duncan, 2001; Lyutikov, 2006), and possibly to FRBs (Bransgrove et al., 2025).
@@ 5.1 Magnetar flares
Magnetic fields of magnetars are highly active and evolve due to MHD effects in the crust, such as Hall drift (Goldreich & Reisenegger, 1992; Pons et al., 2009; Gourgouliatos et al., 2016; Kojima et al., 2021; Igoshev et al., 2021) and plastic flows (Levin & Lyutikov, 2012; Lyutikov, 2015; Lander, 2016; Gourgouliatos & Lander, 2021); and the core due to ambipolar diffusion and superconducting magnetic field evolution (Graber et al., 2015; Passamonti et al., 2017a, b; Igoshev & Hollerbach, 2023; Skiathas & Gourgouliatos, 2024; Igoshev et al., 2025). These effects lead to the formation of multipolar magnetic structures in the lower magnetosphere and may release them to the magnetosphere. Once an arcade forms in the magnetosphere, it is possible through high twist to become unstable (Parfrey et al., 2013; Ntotsikas & Gourgouliatos, 2025) and disconnect from the stellar magnetic field (Lander et al., 2024; Bransgrove et al., 2025). Such configurations evolve dynamically and give rise to a magnetar flare Figure 8. Magnetar flares are rare explosive events that occur in magnetars. A flare can release immerse amounts of energy reaching $10^{46}$erg in the form of X-rays, shining within a fraction of a second. Moreover, flare ejecta expel masses reaching $10^{24}-10^{26}$ g (Gaensler et al., 2005; Granot et al., 2006), which are subsequently powering scaled-down version of kilonovae through r-process radioactive decay (Cehula et al., 2024).
@@
Figure 8: Schematic illustration of plasmoid formation and emission regimes associated with the different families of self-similar solutions. (a) Magnetic reconnection in the magnetar magnetosphere leads to the formation of a magnetically confined plasmoid embedded within the closed field region. (b) Emission properties of plasmoids corresponding to different mass-loading regimes: P-type solutions (overdense) predominantly emit high-energy radiation (X-rays and $\gamma$-rays), whereas Z- and N-type solutions (underdense, magnetically dominated configurations) correspond to low-density plasmoids capable of producing coherent radio emission.
@@
In the model developed here, the relative content of the spherical configuration in mass, thermal pressure, magnetic and electric energy, temperature profile and the maximum Lorentz factor attained are all related through the choices of the parameters $c_{0}$ and $c_{1}$. Assuming that a blob has a radius of 11km ($10^{5}$ cm) and the maximum value of the magnetic flux corresponds to $10^{25}$ G cm2, so that the magnetic field scales with $10^{15}$G, the magnetic and electric field energies $E_{mag}$ and $E_{el}$ range from $10^{44}$erg to $10^{46}$ erg. The pressure work $W$ ranges from $10^{43}$ erg to $10^{45}$ erg and the temperature is at a level of $10^{11}$. We note that the temperature and energetics evolve with time as the plasmoid expands, with the decrease in electromagnetic and thermal energy of the sphere to be related to the work done against the external medium that the sphere is expanding. The mass load $M$ ranges from $10^{24}$ g to $10^{25}$ g and does not evolve with time.
@@ 5.1.1 Giant flare energetics
Here, magnetar giant flares are interpreted as large-scale, relativistic magnetic reconfigurations described by self-similar, axisymmetric MHD solutions with finite mass loading. For all explored solution families (Z, P, and N), the energy budget at early times is dominated by the electromagnetic component, while the kinetic and thermal contributions depend on the mass-loading and the pressure term which are quantified through parameter $c_{1}$. This is a necessary condition for producing a bright, rapidly varying high-energy spike, since a kinetic- or thermal-energy–dominated outflow would radiate inefficiently at X-ray and $\gamma$-ray energies and primarily power delayed, shock-driven emission. The electromagnetic dominance of the early-time evolution is therefore consistent with the observation that the prompt hard X-$\gamma-ray$ spike contains the majority of the radiated energy in magnetar giant flares (e.g. Hurley et al., 2005; Palmer et al., 2005).
An immediate observational implication is that giant flares may exhibit significant diversity in their late-time signatures despite comparable prompt luminosities. Solutions with larger mass loading (P-type) predict enhanced interaction between the ejecta and the surrounding medium, potentially giving rise to detectable radio afterglows or nebular disturbances, as observed following the 2004 flare of SGR 1806–20 (Gaensler et al., 2005). In contrast, low-mass solutions (Z- and N-type) correspond to clean flares, in which little kinetic energy is available to power long-lived afterglow emission, consistent with the weaker radio signatures of SGR 1900+14 (Frail et al., 1999).
@@ 5.1.2 Timescales and light-curve morphology
with $v_{0}$ of order unity. Dissipation is expected to occur preferentially near the outer boundary of the expanding structure, where strong current sheets and pressure gradients develop. This naturally leads to sub millisecond-scale rise times for the initial hard X-$\gamma-ray$ spike, consistent with the observed rise times of Galactic giant flares (Palmer et al., 2005; Hurley et al., 2005).
The model predicts temporally asymmetric light curves, characterized by a very rapid rise phase followed by a longer decay. The rise time reflects localized dissipation at the expanding boundary, whereas the decay traces the subsequent global magnetic reconfiguration of the system. Since these timescales depend primarily on the expansion speed rather than on the total energy release, similar rise times are expected across giant flares of different luminosities. While the presence of strong currents and discontinuities lead to dissipation and current driven instabilities, a detailed study is required to confirm and follow in detail the development of these effects.
@@ 5.1.3 Angular structure and anisotropy
The meridional structure of the solution leads to highly anisotropic profiles with magnetic field, pressure and mass loading to be different on the equator and the poles. Consequently, the emitted radiation is expected to be anisotropic, with the apparent energy of a flare depending on the observer’s viewing angle relative to the magnetic field symmetry axis.
Such anisotropy provides a natural explanation for the wide dispersion in inferred isotropic-equivalent energies among observed giant flares (Hurley et al., 1999), without requiring intrinsic energy releases spanning several orders of magnitude.
@@ 5.1.4 Mass loading and flare diversity
The explored parameter space spans ejecta masses ranging from effectively negligible values (Z-type solutions) up to $\sim 10^{25}$$10^{26}\,\mathrm{g}$. This range is consistent with estimates of the ejecta mass inferred from modelling the radio afterglow of SGR 1806–20 (Gaensler et al., 2005; Granot et al., 2006). Low-mass events, associated with Z- and N-type solutions, correspond to predominantly magnetospheric reconfigurations with minimal baryon contamination. These events are expected to produce sharp, luminous high-energy spikes and weak or undetectable afterglows.
In contrast, P-type solutions involve substantial mass loading, potentially including baryonic material originating from the stellar crust or inner magnetosphere (Thompson & Duncan, 2001; Bransgrove et al., 2025). Such events are expected to exhibit stronger interaction with the ambient medium, leading to detectable radio afterglows or long-lived nebular emission.
5.1.5 Spectral and polarization signatures
As the system expands self-similarly, the magnetic field strength decreases rapidly with radius, implying fast spectral evolution of the emitted radiation. The model predicts an initially hard spectrum during the prompt spike, followed by rapid hard-to-soft evolution on millisecond timescales, in agreement with observed giant flare spectra (Feroci et al., 2001; Olive et al., 2004).
Because the magnetic field remains globally ordered during the early stages of the expansion, the prompt emission is expected to be highly linearly polarized, with a relatively stable polarization angle during the rise phase. This is consistent with expectations for magnetically dominated emission regions (Lyutikov, 2006). Significant depolarization may occur only at later times, once magnetic reconnection and dissipation disrupt the large-scale field structure.
@@ 5.2 Fast radio bursts
Within the present framework, FRBs are interpreted as the coherent radio counterparts of magnetically dominated explosions with low baryon loading Lyubarsky (2021). While the total energy release and large-scale dynamics may resemble those of magnetar giant flares, the conditions required for efficient coherent emission impose strong constraints on the ejecta mass and magnetization. In particular, FRB production is favoured in solutions where the electromagnetic energy dominates over both kinetic and thermal components throughout the early expansion (Lyutikov, 2017; Beloborodov, 2017).
In heavily mass-loaded explosions, such as P-type solutions, the inertia of the ejecta and the rapid formation of a strong compression shock suppress coherent emission by reducing the magnetization and increasing the plasma frequency. These events are therefore expected to produce luminous high-energy flares but weak or absent radio bursts. In contrast, Z- and N-type solutions correspond to minimally mass-loaded outflows in which the plasma remains highly magnetized and optically thin, allowing coherent radio emission to escape. This qualitative distinction is consistent with the observed diversity of magnetar activity, including radio-quiet giant flares and radio-loud but energetically modest bursts (Kaspi & Beloborodov, 2017; Mereghetti et al., 2020).
The self-similar expansion naturally leads to the formation of strong current sheets and charge-starved regions near the expanding boundary of the magnetized structure. In such regions, large-amplitude electromagnetic disturbances can be generated on millisecond timescales, providing favourable conditions for coherent curvature radiation or synchrotron maser emission at relativistic shocks (Lyubarsky, 2014; Plotnikov & Sironi, 2019). The characteristic duration of the resulting radio burst is set by the causal timescale $t\sim r/c$, consistent with the millisecond durations of observed FRBs.
An important prediction of the model is that FRBs need not be associated with the most energetic magnetar flares. Instead, they preferentially accompany low-mass explosions in which a large fraction of the released energy remains in electromagnetic form rather than being degraded into kinetic energy at shocks (Lyutikov, 2022). This provides a natural explanation for the observed association of FRBs with magnetars (Popov & Postnov, 2013), including the detection of a bright radio emission coincident with an X-ray flare from SGR 1935+2154 (Cameron et al., 2005; Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020), as well as the absence of FRBs during the most extreme giant flares.

### v2

@@ 4.2.2 Mass
The density profile of the plasmoid is related to the flux function $\Psi$, Eqs. (17) and (21), multiplied by the prefactor $\Gamma^{3}v^{3}/r^{3}$, as demonstrated in Eq. (15). The two families of solutions we studied here for positive and negative $c_{1}$ have qualitatively different density profiles. For $c_{1}>0$, the minimum is located at the centre of the sphere ($r=0$) and peaks close to the location of the maximum of $\Psi$ with some modulation due to the $\Gamma^{3}$ prefactor that pushes the maximum closer to the surface. In contrast, for $c_{1}<0$, the minimum is located close to the maximum of $\Psi$.
We first considered the rest mass contained within the expanding plasmoid. We integrated the rest-mass density expression given by Eq. (15) within a sphere of given radius $R$, so that $R=cv_{0}\penalty 10000\ t$ at some time $t$, which corresponds to the physical location of the non-zero root of $g$. By accounting for the axial symmetry of the system, integration over $\phi$ provides a factor of $2\pi$,
@@
We confirm, as expected, that the rest mass contained within a spherical configuration does not evolve with time or the physical radius of the fluid, as the integral is on the $v$-space, where the solution does not evolve. However, the amount of mass contained inside the sphere depends on the solution. In what follows, we assume that the function $Q$ is constant and sets the scale for the mass. Furthermore, the total mass is sensitive to the choice of the additive constant, $\tilde{p}_{0}$. We made the following choices: for $c_{1}>0$, we set $\tilde{p}_{0}=0$, whereas for $c_{1}<0$, we chose the minimum acceptable value of $\tilde{p}_{0}$ so that the minimum of the density is exactly zero, and thus, $\tilde{p}_{0}=-\frac{c_{1}}{16\pi^{3}}$. The resulting masses from the integration are shown in Fig. 7. We note that for higher absolute values of $c_{1}$ the mass is higher. We note that the assumption of an additive constant affects the direct comparison of the mass: the P-type solutions have zero density on their surface, where the N-type solutions have a non-zero density on their surfaces. To produce a reasonable comparison, the P-type solutions have a higher mass load than the N-type solutions for the same density on the surface $v=v_{0}$.
Figure 7: Mass $M$ and work $W$ for the P and N solutions as functions of $|c_{1}|$. While $M$ and $W$ are plotted on the same axis, the units differ. Moreover, $W$ evolves with time, which is inversely proportional to $t$.
@@ 4.2.3 Thermal pressure work
Following a similar procedure as for the density, we evaluated the work $W$ of the thermal pressure. It can be calculated through the integral of $pdV$. Upon integration for a sphere of radius $R$ at time $t$, we obtain the following expression:
@@
We note that $W$ is inversely proportional to $t$, signifying the expansion of the system. This work is associated with the spatial redistribution of the plasma mass and pressure as the system expands. We further note that the thermal pressure on the surface of the sphere decreases with time, and thus, as the system expands, less work in done against the external medium. The results of this integration for various choices is shown in Fig. 7 with dotted lines.
@@ 4.2.4 Temperature
Given the relation between the density and the pressure, we evaluated the temperature of the plasmoid through the relation $T=\frac{m\penalty 10000\ p}{k_{B}\rho}$, where $m$ is the particle mass, and $k_{B}$ is the Boltzmann constant. Therefore, the temperature is given by the form:
$T=\frac{m}{k_{B}}\frac{\tilde{p}}{\tilde{\rho}}\frac{\Gamma v}{r}=\frac{m}{k_{B}}\frac{\tilde{p}}{\tilde{\rho}}\frac{\Gamma}{ct}.$ (31)
@@
As $\tilde{p}$ and $\tilde{\rho}$ are related to each other through $Q$, we followed the assumption of a constant value of $Q(v,\theta)$ in the evaluation of the temperature, as we already chose in the calculation of mass through the density. Therefore, we expect a temperature profile for the spherical structure that decreases inversely proportionally with time.
For positive values of $c_{1}$, the maximum value of the temperature approximately coincides with the maximum of the flux function $\Psi$, while it is cooler at the boundaries of the sphere and the axis (Fig. 4 panels c and g). For negative values of $c_{1}$, the minimum of the temperature occurs close to the maximum of $\Psi$, while the surroundings are hotter, as illustrated in Fig. 5 panels c and g. For the normalisation values assumed here, we refer to the caption of Table LABEL:tab:1, and when we assume a proton-electron plasma, the units of the temperature correspond to $10^{11}$K.
@@ 5 Astrophysical implications
The solutions derived here provide a framework for the study of astrophysical systems of relativistically expanding configurations of magnetised plasma. It has been proposed that these structures might arise from magnetars expelling magnetic plasmoids, which are associated with magnetar flares (Thompson and Duncan, 2001; Lyutikov, 2006), and possibly to FRBs (Bransgrove et al., 2025).
@@ 5.1 Magnetar flares
The magnetic fields of magnetars are highly active and evolve due to MHD effects in the crust, such as the Hall drift (Goldreich and Reisenegger, 1992; Pons et al., 2009; Gourgouliatos et al., 2016; Kojima et al., 2021; Igoshev et al., 2021), and plastic flows (Levin and Lyutikov, 2012; Lyutikov, 2015; Lander, 2016; Gourgouliatos and Lander, 2021), and to effects the core due to ambipolar diffusion and to the evolution of the superconducting magnetic field (Graber et al., 2015; Passamonti et al., 2017a, b; Igoshev and Hollerbach, 2023; Skiathas and Gourgouliatos, 2024; Igoshev et al., 2025). These effects lead to the formation of multipolar magnetic structures in the lower magnetosphere and may release them to the magnetosphere. When an arcade forms in the magnetosphere, it might become unstable through high twist (Parfrey et al., 2013; Ntotsikas and Gourgouliatos, 2025) and can disconnect from the stellar magnetic field (Lander et al., 2024; Bransgrove et al., 2025). Such configurations evolve dynamically and give rise to a magnetar flare Fig. 8. Magnetar flares are rare explosive events that occur in magnetars. A flare can release immense amounts of energy that can reach $10^{46}$erg in the form of X-rays, shining within a fraction of a second. Moreover, flare ejecta expel masses reaching $10^{24}-10^{26}$ g (Gaensler et al., 2005; Granot et al., 2006), which subsequently power scaled-down versions of kilonovae through r-process radioactive decay (Cehula et al., 2024).
@@
Figure 8: Schematic illustration of plasmoid formation and emission regimes associated with the different families of self-similar solutions. (a) Magnetic reconnection in the magnetar magnetosphere leads to the formation of a magnetically confined plasmoid embedded within the closed-field region. (b) Emission properties of plasmoids corresponding to different plasmoid regimes: P-type solutions (overdense) predominantly emit high-energy radiation (X-rays and $\gamma$-rays), whereas Z- and N-type solutions (underdense magnetically dominated configurations) correspond to low-density plasmoids capable of producing coherent radio emission.
@@
In the model developed here, the relative content of the spherical configuration in mass, thermal pressure, magnetic and electric energy, temperature profile, and the maximum Lorentz factor attained are all related through the choices of the parameters $c_{0}$ and $c_{1}$. When we assume that a blob has a radius of 11km ($10^{5}$ cm) and the maximum value of the magnetic flux corresponds to $10^{25}$ G cm2, so that the magnetic field scales with $10^{15}$G, the magnetic and electric field energies $E_{mag}$ and $E_{el}$ range from $10^{44}$erg to $10^{46}$ erg. The pressure work $W$ ranges from $10^{43}$ erg to $10^{45}$ erg, and the temperature is at a level of $10^{11}$. We note that the temperature and energetics evolve with time as the plasmoid expands, with the decrease in electromagnetic and thermal energy of the sphere related to the work done against the external medium into which the sphere is expanding. The mass load $M$ ranges from $10^{24}$ g to $10^{25}$ g and does not evolve with time.
@@ 5.1.1 Giant flare energetics
We interpret magnetar giant flares as large-scale relativistic magnetic reconfigurations described by self-similar axisymmetric MHD solutions with finite mass loading. For all explored solution families (Z, P, and N), the energy budget at early times is dominated by the electromagnetic component, while the kinetic and thermal contributions depend on the mass loading and the pressure term, which are quantified through parameter $c_{1}$. This is a necessary condition for producing a bright, rapidly varying high-energy spike because an outflow dominated by kinetic or thermal energy would radiate inefficiently at X-ray and $\gamma$-ray energies and primarily power delayed shock-driven emission. The electromagnetic dominance of the early-time evolution is therefore consistent with the observation that the prompt hard X-ray or $\gamma$-ray spike contains the majority of the radiated energy in magnetar giant flares (e.g. Hurley et al., 2005; Palmer et al., 2005).
An immediate observational implication is that giant flares can exhibit significant diversity in their late-time signatures despite comparable prompt luminosities. Solutions with higher mass loading (P-type) predict enhanced interaction between the ejecta and the surrounding medium, potentially giving rise to detectable radio afterglows or nebular disturbances, as observed following the 2004 flare of SGR 1806–20 (Gaensler et al., 2005). In contrast, low-mass solutions (Z- and N-type) correspond to clean flares, in which little kinetic energy is available to power long-lived afterglow emission, consistent with the weaker radio signatures of SGR 1900+14 (Frail et al., 1999).
@@ 5.1.2 Timescales and light-curve morphology
with $v_{0}$ of order unity. Dissipation is expected to occur preferentially near the outer boundary of the expanding structure, where strong current sheets and pressure gradients develop. This naturally leads to sub-millisecond rise times for the initial hard X-ray or $\gamma$-ray spike, consistent with the observed rise times of Galactic giant flares (Palmer et al., 2005; Hurley et al., 2005).
The model predicts temporally asymmetric light curves, characterised by a very rapid rise phase followed by a longer decay. The rise time reflects localised dissipation at the expanding boundary, whereas the decay traces the subsequent global magnetic reconfiguration of the system. Since these timescales primarily depend on the expansion speed and not on the total energy release, similar rise times are expected across giant flares of different luminosities. While the presence of strong currents and discontinuities lead to dissipation and current-driven instabilities, a detailed study is required to confirm and follow the development of these effects in detail.
@@ 5.1.3 Angular structure and anisotropy
The meridional structure of the solution leads to highly anisotropic profiles with magnetic field, pressure, and mass loading different on the equator and at the poles. Consequently, the emitted radiation is expected to be anisotropic, with the apparent energy of a flare depending on the observer’s viewing angle relative to the magnetic field symmetry axis.
This anisotropy provides a natural explanation for the wide dispersion in inferred isotropic-equivalent energies among observed giant flares (Hurley et al., 1999), without requiring intrinsic energy releases spanning several orders of magnitude.
@@ 5.1.4 Mass loading and flare diversity
The explored parameter space spans ejecta masses ranging from effectively negligible values (Z-type solutions) to $\sim 10^{25}$$10^{26}\,\mathrm{g}$. This range is consistent with estimates of the ejecta mass inferred from modelling the radio afterglow of SGR 1806–20 (Gaensler et al., 2005; Granot et al., 2006). Low-mass events, associated with Z- and N-type solutions, correspond to predominantly magnetospheric reconfigurations with minimum baryon contamination. These events are expected to produce sharp, luminous high-energy spikes and weak or undetectable afterglows.
In contrast, P-type solutions involve substantial mass loading, potentially including baryonic material originating from the stellar crust or inner magnetosphere (Thompson and Duncan, 2001; Bransgrove et al., 2025). Such events are expected to exhibit stronger interaction with the ambient medium, leading to detectable radio afterglows or long-lived nebular emission.
5.1.5 Spectral and polarisation signatures
As the system expands self-similarly, the magnetic field strength decreases rapidly with radius, implying a fast spectral evolution of the emitted radiation. The model predicts an initially hard spectrum during the prompt spike, followed by rapid hard-to-soft evolution on millisecond timescales, in agreement with observed giant flare spectra (Feroci et al., 2001; Olive et al., 2004).
Because the magnetic field remains globally ordered during the early stages of the expansion, the prompt emission is expected to be highly linearly polarised, with a relatively stable polarisation angle during the rise phase. This is consistent with expectations for magnetically dominated emission regions (Lyutikov, 2006). Significant depolarisation might only occur at later times, after magnetic reconnection and dissipation disrupted the large-scale field structure.
@@ 5.2 Fast radio bursts
Within the present framework, FRBs are interpreted as the coherent radio counterparts of magnetically dominated explosions with low baryon loading Lyubarsky (2021). While the total energy release and large-scale dynamics can resemble those of magnetar giant flares, the conditions required for efficient coherent emission impose strong constraints on the ejecta mass and magnetisation. In particular, FRB production is favoured in solutions where the electromagnetic energy dominates the kinetic and thermal components throughout the early expansion (Lyutikov, 2017; Beloborodov, 2017).
In heavily mass-loaded explosions, such as P-type solutions, the inertia of the ejecta and the rapid formation of a strong compression shock suppress coherent emission by reducing the magnetisation and increasing the plasma frequency. These events are therefore expected to produce luminous high-energy flares, but weak or absent radio bursts. In contrast, Z- and N-type solutions correspond to minimally mass-loaded outflows in which the plasma remains highly magnetised and optically thin, allowing coherent radio emission to escape. This qualitative distinction is consistent with the observed diversity of magnetar activity, including radio-quiet giant flares and radio-loud but energetically modest bursts (Kaspi and Beloborodov, 2017; Mereghetti et al., 2020).
The self-similar expansion naturally leads to the formation of strong current sheets and charge-starved regions near the expanding boundary of the magnetised structure. In these regions, large-amplitude electromagnetic disturbances can be generated on millisecond timescales, providing favourable conditions for coherent curvature radiation or synchrotron maser emission at relativistic shocks (Lyubarsky, 2014; Plotnikov and Sironi, 2019). The characteristic duration of the resulting radio burst is set by the causal timescale $t\sim r/c$, which is consistent with the millisecond durations of observed FRBs.
An important prediction of the model is that FRBs need not be associated with the most energetic magnetar flares. Instead, they preferentially accompany low-mass explosions in which a large fraction of the released energy remains in electromagnetic form and is not degraded into kinetic energy at shocks (Lyutikov, 2022). This provides a natural explanation for the observed association of FRBs with magnetars (Popov and Postnov, 2013), including the detection of a bright radio emission coincident with an X-ray flare from SGR 1935+2154 (Cameron et al., 2005; Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020), and the absence of FRBs during the most extreme giant flares.